### v1

# X-ray panorama of the SS 433/W50 complex by SRG/eROSITA

X-ray panorama of the SS 433/W50 complex by SRG/eROSITA
Galactic microquasar SS 433 and the radio nebula W50 surrounding it present a prototypical example of a hyper-Eddington binary system shaping its ambient interstellar medium via energetic outflows. In this paper, we present X-ray observations of the SS 433/W50 complex by the eROSITA telescope onboard the SRG space observatory. These data provide images of the entire nebula characterized by a very large dynamic range and allow spectral analysis of the diffuse X-ray emission. In particular, these data illustrate a close connection between the thermal and non-thermal components of W50 on scales ranging from sub-parsecs, represented by narrow X-ray bright filaments, to the entire extent of the nebula. These data also allow us to fully characterize a pair of nearly symmetric, sharp-edged, elongated structures aligned with the orbital axis of the binary system, which lack radio counterparts, but are prominent in very high energy gamma-ray emission. The resulting multifaceted picture of the interaction between energetic outflows and the surrounding medium paves the way for future focused multiwavelength observations and dedicated numerical simulations.
Key Words.:
ISM: jets and outflows – X-rays: binaries – plasmas – acceleration of particles1 Introduction
SS 433 is a unique close stellar binary system in our Galaxy (located at and kpc) composed of a relativistic compact object (a black hole or a neutron star) which accretes matter from its companion at sustainably hyper-Eddington rate (see Fabrika, 2004; Cherepashchuk et al., 2025, for reviews). Having been first listed in the catalog of point-like H emitters (Stephenson & Sanduleak, 1977), it quickly became one of the most studied astrophysical objects thanks to its very peculiar optical spectrum (Margon et al., 1979) featuring pairs of narrow (with half-opening angle ) blue- and red-shifted lines (with respect to their rest frame positions) with the shifts varying periodically in time (see Margon, 1984, for a review of early observational results). A kinematic model associates this emission with a pair of narrow relativistic jets having constant bulk velocity of a quarter of the speed of light and changing direction in a regular precession manner with a period days (Abell & Margon, 1979; Fabian & Rees, 1979). This period is much longer than the period of orbital modulation ( days) and is interpreted as due to precession of a thick accretion disk with the amplitude of 21∘ and the mean axis inclined by the angle to the line of sight (e.g., Katz, 1980; Sarazin et al., 1980; Whitmire & Matese, 1980). Subsequent detection of extended and variable radio emission (Spencer, 1979) perfectly described by the precessing jets pattern fully confirmed the picture of twin relativistic outflows (Hjellming & Johnston, 1981a, b; Blundell & Bowler, 2004). Moreover, the proper motion of the individual radio-emitting blobs allowed one to constrain the distance to the source with a remarkable 10% accuracy, being kpc (Hjellming & Johnston, 1981a; Blundell & Bowler, 2004; Marshall et al., 2013).
Despite indications of both quasi-regular (nutation, Katz et al., 1982) and sporadic (jittering, Kubota et al., 2010) deviations from the predictions of this simple model as well as presence of (relatively short) periods of time when the Doppler shifts of optical emission lines change significantly (e.g., Blundell et al., 2011; Medvedev et al., 2022), all the derived parameters of the model stay remarkably stable over more than 40 years of continuous monitoring observations (Cherepashchuk et al., 2018). This stability likely reflects extreme robustness of the mass transfer occurring in the hyper-Eddington regime (van den Heuvel, 1981; van den Heuvel et al., 2017) with the estimated accretion rate (e.g., Shklovskii, 1981; Cherepashchuk, 1981; Leibowitz, 1984; Fuchs et al., 2006), where is the mass of the compact object and is the Solar mass and is the critical Eddington mass accretion rate. The compact object is most likely a black hole with (Cherepashchuk et al., 2023), although the possibility of it being a neutron star () has also been discussed (e.g. Goranskij, 2011; Medvedev et al., 2018).
Since X-ray emission coming from the innermost parts of the hyper-Eddington accretion disk is expected to be strongly beamed in the axial direction (Shakura & Sunyaev, 1973; Begelman et al., 2006; Poutanen et al., 2007), we are not able to observe it directly, given the high inclination angle of the system. Indeed, although X-ray emission from this source has also been detected early on (Seward et al., 1976; Forman et al., 1978), its luminosity is relatively low compared to the Eddington luminosity (Grindlay et al., 1984). As a result, SS 433 has long been considered as a beamed away prototype of Ultra Luminous X-ray sources (ULXs, e.g., Kaaret et al., 2017, for a review) observed in normal star-forming galaxies (Fabrika & Mescheryakov, 2001; Begelman et al., 2006; Poutanen et al., 2007), although possible indications of the ‘hidden’ central emission are yet to be confirmed (Medvedev & Fabrika, 2010; Khabibullin & Sazonov, 2016, 2019; Middleton et al., 2021; Fogantini et al., 2023).
Instead, what dominates X-ray emission of the system is again thermal emission from a twin pair of the baryonic jets having almost identical orientation and bulk velocity to the optically emitting jets (Watson et al., 1986). Numerous pairs of blue- and red-shifted narrow emission lines of highly ionized heavy elements (primarily, silicon, sulfur, iron and nickel) allow one to build a multi-temperature model of the ballistically expanding and cooling flow (Watson et al., 1986; Kotani et al., 1996; Marshall et al., 2002; Brinkmann et al., 2005; Medvedev & Fabrika, 2010; Marshall et al., 2013; Khabibullin et al., 2016) which first becomes visible when its temperature is above 20 keV and then cools down below 1 keV and likely fragments later on due to thermal instability (Brinkmann et al., 1988). Detailed spectroscopic modeling of this emission (e.g. Khabibullin et al., 2016; Medvedev et al., 2019) in combination with the eclipsing observations of the jets by the companion star (Stewart et al., 1987; Kawai et al., 1989; Filippova et al., 2006; Lopez et al., 2006; Atapin & Fabrika, 2016) enable very robust determination of the jets’ physical properties (so-called jets ‘biometry’, consisting of their size, opening angle, velocity, temperature, and the mass flux). Thanks to this we know very robustly the kinetic luminosity of these outflows, which exceeds (e.g., Medvedev et al., 2019), confirming the necessity of a hyper-Eddington ‘engine’ releasing large amount of energy in the form of relativistic outflows, consistently with the early theoretical (Shakura & Sunyaev, 1973) and modern numerical (e.g., Ohsuga & Mineshige, 2011; Jiang et al., 2019; Yoshioka et al., 2022; Fragile et al., 2025) predictions.
In contrast to the radiation freely escaping the system, the kinetic energy of the launched outflows cannot be easily transported away from the source and has to be deposited into the surrounding interstellar medium. The estimated total amount of this energy exceeds erg for a life-time of the system at the level of 30,000 yr (e.g., van den Heuvel et al., 1980). This energy is at least on par with the energy of the initial supernova explosion that should have accompanied the birth of the compact object, meaning that a quite large region pc in its vicinity can be shaped by this later activity (e.g., Begelman et al., 1980; Zealey et al., 1980; van den Heuvel, 1981; Velázquez & Raga, 2000; Zavala et al., 2008; Goodall et al., 2011a; Churazov et al., 2020, 2024).
Indeed, SS 433 is known to be surrounded by the famous giant ( pc) radio nebula W50 (Westerhout, 1958; Holden & Caswell, 1969; Geldzahler et al., 1980; Elston & Baum, 1987; Dubner et al., 1998; Goodall et al., 2011b; Broderick et al., 2018; Sakemi et al., 2021), which has a suggestive elongated morphology, with the elongation axis being co-aligned with the symmetry axis of the jets precession cone. The asymmetry of the lobes, or “ears”, of the nebula can be readily explained by the noticeable stratification of the surrounding medium, given that the system is located at the height pc below the Galactic plane and has the size comparable to the scale height of the exponential cold gas disk (e.g., Velázquez & Raga, 2000; Zavala et al., 2008; Goodall et al., 2011a). The interior of the nebula is also known to shine in soft thermal X-rays (Brinkmann et al., 1996, 2007; Safi-Harb & Ögelman, 1997; Chi et al., 2024), hard non-thermal (Seward et al., 1980; Watson et al., 1983; Yamauchi et al., 1994; Safi-Harb & Petre, 1999; Namiki et al., 2000; Safi-Harb et al., 2022; Kaaret et al., 2024) and filamentary H emission (Kirshner & Chevalier, 1980; Shuder et al., 1980; Boumis et al., 2007; Abolmasov et al., 2010; Farnes et al., 2017; Rosado et al., 2021), indicative of the high velocity shocks propagating through a relatively dense interstellar medium.
The nature of the extended X-ray jets (EXJ) remains more puzzling, although their morphology, spectra, and polarization all point to the synchrotron origin of the emission (e.g., Safi-Harb et al., 2022; Kaaret et al., 2024). Most recently, Very High Energy (VHE) emission from these structures has been detected and mapped (e.g., Abeysekara et al., 2018; Alfaro et al., 2024; H. E. S. S. Collaboration et al., 2024; LHAASO Collaboration, 2024), demonstrating that indeed ultra-relativistic electrons with energy above 100 TeV are accelerated in the system (e.g, Kimura et al., 2020), probably as a consequence of internal shocks and recollimation in the energetic trans-relativistic outflows (Bykov et al., 2025). The overall production of cosmic rays by such systems (super-Eddington accretors with powerful outflows) might be of importance for the total CR budget and local variations in its spectrum from place to place (e.g., Heinz & Sunyaev, 2002), in particular in the PeV regime (e.g., Peretti et al., 2025; Wang et al., 2025; Bykov et al., 2025).
Although the paradigm of “a super-Eddington X-ray binary powering a nebula with energetic outflows” appears to be well established and consistent with the most salient observational properties of the SS 433/W50 complex, the details of almost all aspects of this interaction remain unclear. Indeed, all the phenomena directly associated with the central source can be trace only up to a distance of mere 0.1 pc from it, where the last signatures of the corkscrew pattern of the precessing jets are observed in radio and X-rays (Migliari et al., 2005; Miller-Jones et al., 2008; Khabibullin & Sazonov, 2017; Sakai et al., 2025). A model invoking steady deceleration and disruption of the jets can, in principle, account for such behavior (Panferov, 2014, 2017), but then the re-appearance of the “jets” with similar to the injection velocities tens of pc away from the central source might be problematic. Morphology of the soft X-ray emission, best mapped by the ROSAT mosaic observations (Brinkmann et al., 1996, 2005), is also quite puzzling, with large global brightness variations and thin filamentary structures, not very commonly observed in similar situations of supernova remnants (which typically feature shell-, plerion- or mixed types of morphology). Deep pointed observations with XMM-Newton, Chandra, Suzaku, and NuSTAR were capable of covering certain parts of the nebula (e.g., Safi-Harb et al., 2022; Chi et al., 2024; Tsuji et al., 2025), but a uniform and sensitive coverage of the entire nebula with adequate spectroscopic capabilities is missing.
Here we present results of X-ray observations of the SS 433/W50 complex by the eROSITA telescope (Predehl et al., 2021) onboard the SRG space observatory (Sunyaev et al., 2021), which were performed during its Performance Verification (PV) phase. The obtained mosaic provides us with X-ray images of the entire extent of W50 with unprecedented dynamic range, bringing X-ray data on par with the radio and other multi-wavelength data, and, in addition, delivers exquisite spectral information for the rich variety of diffuse structures. These data, in particular, demonstrate a close connection between the thermal and non-thermal components of W50 on scales ranging from subparsec, represented by narrow X-ray and radio bright filaments, to the entire extent of the nebula. These data also enable full characterization of a pair of nearly symmetric, sharp-edged, elongated structures aligned with the orbital axis of the binary system, which lack radio counterparts.
The structure of the paper is following: we describe the observations and the obtained data in Sect. 2; results of the broad-band and spectrally-resolved imaging are presented in Sect. 3 and Sect. 4, respectively; spectroscopic analysis of the emission from several selected regions follows in Sect. 5; the obtained global picture is discussed in Sect. 6 and summarized in Sect. 7.
2 Data
Since W50 spans approximately degrees on the sky, corresponding to the physical size of at the distance of 5 kpc111The most recent parallax measurement for SS 433 in Gaia DR3 corresponds to the significantly larger distance of kpc (Gaia DR3 4293406612283985024, Gaia Collaboration et al., 2023), we caution of using this value given large residuals of the astrometric solution. The previous parallax measurement published in Gaia DR2 corresponded to a much smaller distance of kpc (Gaia DR2 4293406612283985024, Arnason et al., 2021)., the observing campaign consisted of 9 individual telescope pointings (each covering a circular area of in diameter). These pointings were arranged in a mosaic to result in a nearly uniform coverage of the entire nebula and, in addition, the highest sensitivity in the most interesting regions. The observing dates, from October 5th to October 8th and then from October 19th to October 20th, 2019, were chosen to coincide with the two consecutive orbital eclipses of the central source SS 433 by its normal companion star (according to orbital ephemerids by Eikenberry et al., 2001). During eclipses, the apparent X-ray flux of SS 433 goes down by a factor of several (e.g., Kawai et al., 1989; Lopez et al., 2006; Medvedev et al., 2019), minimizing possible adverse effects of contamination and pile-up due to the presence of a very bright object in the field-of-view, even though relative suppression of the soft X-ray lines being rather minor (Marshall et al., 2013). The typical effective (i.e., equivalent to the nominal eROSITA on-axis performance) exposure time across the map amounts to 30-50 ks, leading to a factor of deeper images compared to the previously available full-extent maps of W50 by observatory in the 0.5-2 keV band (Brinkmann et al., 1996) and Einstein observatory in the 2-4 keV band (Watson et al., 1983). In addition to that, the eROSITA data have much better spatial ( arcsec on the telescope axis and averaged across the field-of-view) and spectral ( eV at 1 keV) resolution (Predehl et al., 2021). A comparable quality of the data is provided by multiple Chandra and XMM-Newton observations, but their coverage of the whole W50 nebula is non-uniform and still incomplete after more than 20 years of observations (Brinkmann et al., 2005; Safi-Harb et al., 2022; Chi et al., 2024). On the other hand, the SRG/eROSITA data are complementary to these observations, as well as to the imaging observations in harder X-rays with NuSTAR, which offer a unique view of non-thermally emitting populations. In this paper, we focus on the SRG/eROSITA X-ray data alone and leave combined analyses for future studies.
We also supplement the PV data with the data obtained in the course of the SRG/eROSITA All-Sky Survey (Predehl et al., 2021; Sunyaev et al., 2021). Although the survey data are much shallower in sensitivity (the typical exposure time per point is s after combining the data of four scans), they provide uniform coverage of the entire area surrounding the nebula. For them, the standard data reduction, background estimation and exposure- and vignetting corrections were applied in the same way as was done for mapping diffuse emission from nearby galaxy clusters (Churazov et al., 2021a, 2023), supernova remnants (Churazov et al., 2021b; Khabibullin et al., 2023, 2024), and the Galactic diffuse emission (Predehl et al., 2020; Khabibullin et al., 2022; Churazov et al., 2024) on even larger scales. Here we also go one step further and combine the PV and all-sky data in a seamless fashion, similarly as it has been done for Chandra and SRG/eROSITA observations of the Perseus cluster (Churazov et al., 2025).
A comparison of the degrees field in radio and X-ray bands is shown in Fig. 1. Although in the radio image, the outer boundary of the W50 nebula appears to be very prominent, the X-ray image is instead dominated by the very bright central object (SS 433) and elongated structures known as “extended X-ray jets” (EXJ) located well inside the radio boundary of W50 (Seward et al., 1980; Watson et al., 1983; Brinkmann et al., 1996). At a distance of 5 kpc, these images span 350 pc and range from 17 pc to almost 370 pc in distance from the Galactic plane. Given that the scale heights for the cold and warm components of the interstellar medium are tens and hundreds of pc, respectively, one could naturally expect signatures of the environment stratification (e.g., Goodall et al., 2011a). This is indeed revealed by the asymmetry of the W50 lobes, with the one directed towards the Galactic plane (the Western Lobe in equatorial coordinates) being more compact than the opposite one (the Eastern lobe).
On larger scales, clear variations of the background and foreground diffuse X-ray emission are seen. Inspection of 3D reddening maps (e.g., Green et al., 2019) shows a clear anti-correlation of the reddening and soft X-ray surface brightness (see Appendix C), suggesting that absorption of X-rays is the main reason for these variations. Morphologically similar structures appear in the 3D reddening maps at a distance of 2-3 kpc, i.e., in the foreground to W50/SS 433. Therefore, it is plausible that the majority of them are physically unrelated to W50.
3 Broad-band imaging
We now turn to the deeper PV data. The broad-band (0.5-4 keV) vignetting-corrected X-ray image of the entire W50 nebula is shown in Fig. 2, with the contours of radio emission at 4.85 GHz (Gregory et al., 1996) overlaid in white. The color coding reflects the measured X-ray surface brightness on a square-root scale spanning two orders of magnitude, highlighting the wide dynamic range of the diffuse X-ray emission captured by eROSITA. In addition to SS 433 itself (the extremely bright source in the center) and a multitude of point sources, the majority of which are nearby foreground stars or background AGN, the picture shows bright diffuse emission with rich internal structure.
The X-ray picture can be broadly decomposed into three components. Firstly, there is a low surface brightness emission filling the entire projected extent of the nebula. It is brighter in the Eastern and Western lobes of W50 (left and right, respectively, in Fig. 2) and significantly dimmer along the Northern and Southern radio boundaries of W50. Remarkably, the central part of W50 (beyond the region contaminated by the PSF wings from SS 433) appears even fainter in X-rays, suggesting a shell-like geometry of the X-ray-emitting gas.
Secondly, one can notice the presence of very narrow X-ray filaments, particularly inside and close to the boundaries of the Eastern and Western lobes. Their apparent width is , while their length is . Given the angular resolution of eROSITA (better than , i.e., at 5 kpc), these filaments are fully resolved in both directions. The X-ray filaments form a complicated network with an overall structure very similar to the filaments of the cold and dense gas observed via H+[N II] emission at different locations within the nebula (Boumis et al., 2007; Farnes et al., 2017). The largest X-ray filaments have a flat transverse profile of the surface brightness and are an order of magnitude brighter than the surrounding medium. Such filaments of the dense X-ray emitting gas might arise as instabilities in the sheared flows, caustics of the strong shock waves, or due to heating of the preexisting cold gas filaments by hot gas and relativistic particles inside the lobes.
The final and perhaps the most visual feature of the image is a pair of nearly symmetrical, sharp-edged “extended X-ray jets” directed away from the central source. Their axis is not only aligned with the elongation of the radio nebula and location of the X-ray bright lobes, but also coincides with the projection of the precession axis of SS 433’s famous trans-relativistic baryonic jets. The physical nature of these structures, first found by the Einstein observatory (Watson et al., 1983) and then observed with all major X-ray telescopes (Brinkmann et al., 1996; Safi-Harb & Ögelman, 1997; Namiki et al., 2000; Brinkmann et al., 2007; Safi-Harb & Petre, 1999), including the most recent polarimteric observations with IXPE (Kaaret et al., 2024), remains a matter of debate (e.g. Churazov et al., 2024). The SRG/eROSITA data clearly show a very rapid onset of emission at the innermost boundary of EXJs on both sides of the nebula (also demonstrated recently by Tsuji et al., 2025, using Chandra data).
4 Spectrally-decomposed imaging
We take advantage of the unique possibility provided by /eROSITA to probe the spectral characteristics of diffuse X-ray emission throughout the full extent of the nebula for the first time. To produce a spectrally-decomposed image, we divided the full spectral band presented earlier into three sub-bands: soft (0.5-1 keV), medium (1-2 keV), and hard (2-4 keV) bands, respectively. A composite RGB image showing the surface brightness of the emission (now on a linear scale) in these bands in red, green, and blue, respectively, is presented in Figure 3. The chosen spectral decomposition immediately reveals the presence of two distinct components in the observed diffuse emission: large regions of red-orange color where emission below 1 keV dominates, and green-cyan structures where prominent emission above 1 keV is present (including the bright central source SS 433). Compact sources also form two groups: the softer (“red-yellow”) ones are predominantly foreground stars, while harder (“blueish”) sources are mostly background AGNs, which shine through the gas and dust in the disk of the Milky Way. Zoom-in images of the lobes (rotated and put side by side to help the comparison) are shown in Fig. 4 and demonstrate fine structure of the soft diffuse emission, contrasting with relatively smooth and well-confined hard emission.
For better characterization of the diffuse emission, we identified bright point sources and masked them along with the central source SS 433. The resulting composite map of the residual diffuse emission is shown in Figure 5. The overlaid distance and angle rulers highlight the characteristic dimensions and morphology of the prominent structures. Namely, one can see that the “Extended X-ray Jets” start pc away from the central source and demonstrate remarkable symmetry, with the Eastern one starting only a couple of parsecs further from the central source. Both EXJs, characterized by hard X-ray spectra, terminate around pc. This symmetry contrasts with the more asymmetric and irregular soft emission of the lobes.
The same conclusion is even better illustrated by the side-by-side comparison of the two lobes in Fig. 4. Indeed, while the Eastern lobe extends up to , the Western one terminates at pc. The inclination of the relativistic jet’s precession axis to the line of sight is (the relativistic jet on the eastern side is on average approaching us), so this difference cannot be explained by geometrical effects only. However, it broadly agrees with the assumption that the energy flow from the central source interacts with the denser ambient medium in the western part, reflecting the gas density gradient perpendicular to the equatorial plane of the Galactic disk (e.g., Goodall et al., 2011a). Since the difference between the lobes’ dimensions amounts to a sizeable fraction of them, one can infer the exponential scale-height of the ambient gas density profile pc, consistent with the scale-height of the Galactic neutral gas disk. The X-ray emission from the Northern and Southern boundaries of the nebula appears soft and faint, being noticeably affected by the foreground dust absorption, especially in their Western parts (as illustrated in Appendix C).
Taking this into account, we conclude that the soft X-ray emission can be traced across the full extent of the nebula, except for the “inner cavity” with a radius of pc. The outer boundary of the emission of the Northern and Southern arcs (outlined as dashed annuli including Region 4 in Fig. 5) is located at from the center, coinciding with the boundary of the W50’s radio emission. This morphology suggests that the soft X-ray emitting gas is confined in a thick () close-to-spherical shell, driving a strong shock in the surrounding ISM.
The Eastern and the Western lobes are roughly confined inside the projection of the cone with the half-opening angle , while the EXJs encompass only the cone with a twice smaller opening angle,. From the eROSITA spectrally decomposed RGB images, a gentle spectral evolution of the EXJs emission with distance is seen clearly. EXJ emission is significantly harder at its sharp inner boundary than further away from SS 433, where EXJs blend with softer thermal emission that fills the rest of the lobe, in agreement with earlier findings with XMM-Newton and NuSTAR (Brinkmann et al., 2007; Safi-Harb et al., 2022).
5 X-ray spectra
5.1 Central source
The spectral capabilities of eROSITA have been demonstrated using many objects, including SS 433 itself (e.g., Sunyaev et al., 2021). Since this source is extremely bright, the observations were intentionally performed during the two consecutive orbital eclipses of the compact object (and the jets) by the companion star, so its soft X-ray flux was dimmer by a factor of several compared to the typical levels (Kawai et al., 1989; Lopez et al., 2006; Marshall et al., 2013). The combined X-ray spectrum of SS 433 extracted from arcmin aperture (with background estimated from to 5 arcmin annuli, being however completely negligible compared the source) using the data of all pointings in which it falls within Field-of-View is shown in Fig. 6.
As expected, it is well described by the model of the optically thin multitemperature emission from collimated flows that adiabatically expand and cool at distances from a few to a few cm from the central object (Brinkmann et al., 1988, 1991, 2005; Kotani et al., 1996; Marshall et al., 2002; Medvedev & Fabrika, 2010; Khabibullin et al., 2016; Medvedev et al., 2019). We discuss it more in Appendix A, while the main focus of this study is instead on the much fainter diffuse emission from the surrounding nebula.
5.2 Diffuse emission
The high sensitivity of the presented data allowed us to perform spectroscopy of the diffuse emission in several regions and determine the physical state of the radiating plasma. We have selected eight regions, shown as numbered boxes in Figure 5, for detailed spectral analysis of their X-ray emission. For each region, the background (estimated from the adjacent regions) was subtracted to isolate the net signal. We present the parameters of the fitted models in Table 1 and discuss the implications of the corresponding results below.
5.2.1 Eastern Lobe
Emission from the brightest filament in the Eastern lobe (Region 1 in Figure 5) features a spectrum characteristic for a thermal (but non-equilibrium) optically-thin plasma with a temperature of , as evidenced by the line emission from the highly ionized oxygen, neon, magnesium, silicon and iron (see the left panel in Figure 7). The line emissivity and ratios are consistent with the Solar abundance of heavy elements, but also indicate a non-equilibrium state of the gas ionization. Namely, it is qualitatively consistent with the expectation for a plane-parallel shell of the gas that experienced a passage of the shock with velocity (from the Rankine-Hugoniot relations for the temperature and density jump properties) some yr ago.
The inferred hydrogen number density, cm-3, is a factor of a few higher than the expected gas density of the interstellar gas pc away from the plane of the Galactic disk. The spectrum of the diffuse emission in the Eastern lobe (exemplified by Region 3 in Figure 5) also can be well described by thermal non-equilibrium emission of the gas with somewhat higher temperature keV and lower hydrogen number density cm-3 (see the left panel in Figure 7).
5.2.2 Western Lobe
The emission from the Western lobe appears to be noticeably harder, partially due to the higher interstellar absorption in that direction. However, it also corresponds to the emission of plasma with genuinely higher temperature, keV, and a similar hydrogen number density, cm-3 (and a few times higher for the brightest filament), as illustrated by spectra from Regions 6 and 7 in Figure 5). The non-equilibrium ionization model (see the right panel in Figure 7) results in a comparable estimate of yr for the ionization time-scale.
The total absorption-corrected luminosity of the diffuse soft X-ray emission from the nebula is in 0.5-4 keV, which is by a factor of few tens smaller than the apparent X-ray luminosity of the central object, few ,222Due to the beaming of the radiation along the thick accretion disc axis (expected in the case of highly supercritical accretion), the actual X-ray luminosity of SS 433 might be orders of magnitude larger, namely comparable to the Eddington luminosity and kinetic luminosity of the jets, (e.g. Fabrika & Mescheryakov, 2001; Begelman et al., 2006; Poutanen et al., 2007; Medvedev & Fabrika, 2010; Khabibullin & Sazonov, 2016; Middleton et al., 2021; Fogantini et al., 2023). and much smaller than the supposed kinetic luminosity of the outflows from the supercritical accretion disk, . However, the thermal energy content of the X-ray emitting gas in the Eastern lobe (with estimated gas number density 0.1 cm-3, temperature 0.5 keV and volume assuming roughly to cylindrical geometry ) is erg, requiring a supply of at least over years. The mass of this gas amounts to , implying that the bulk of it cannot be provided by the outflows from SS 433. If the volume filling factor of the emitting medium is, however, low, the estimate for the required mass and energy could be decreased by a factor of , but similarly, the density should be increased, leading to the corresponding decrease in the time since the shock passage. That means that the X-ray emitting gas provides an estimate of the average energy input in the lobes at the level of , which is comparable to the kinetic luminosity of the jets and the Eddington luminosity of the system (for the mass of the compact object below 10).
A similar estimate can be obtained from the Western lobe, implying nearly symmetric energy injection along the nebula’s axis at a very high rate over the last few tens of thousands of years. Given that the temperature inferred from the spectral fits with the non-equilbirum emission is driven mostly by characteristics of the ionization state, it is likely a lower limit on the actual temperature of the emitting gas. Hence, a significant portion of the gas energy content might be effectively hidden, while the required energy input should be increased correspondingly.
5.2.3 Northern and southern boundaries
The emission from the nebula boundaries is exemplified by the Northern Arc region and the Southern Extension (Regions 4 and 8 in Figure 5). The thermal non-equilibrium model provides the temperature estimate keV, low hydrogen number density, cm-3 for the Northern Arc and even lower for the Southern Extension. The ionization timescale is, however, similar here as well, yr. This observation contrasts with the brightness of the radio emission, which in the Northern Arc region is one of the brightest across the nebula. The high degree of linear polarization of the radio emission might indicate that the preferential orientation of the interstellar magnetic field in this region favors efficient acceleration of relativistic particles producing synchrotron emission. We note here that the central part of the W50 X-ray map suffers from higher photoelectric absorption than the Eastern and even Western lobes, as illustrated in Fig. 9. This excess absorption is likely associated with the foreground gas and has no physical connection to SS 433/W50.
5.2.4 Extended X-ray jets
In contrast to this “thermal” picture, the X-ray spectra of X-ray emission from the “extended X-ray jets” demonstrate a featureless spectrum that can be described by a power law with the slope in the range from (at their base, e.g., Region 5 in Fig. 5) to (at the termination points, e.g. Region 2 in Fig. 5). In agreement with previous spectral studies of this emission (e.g., Brinkmann et al., 2007), the thermal bremsstrahlung model also provides an equally good fit with temperatures from keV (at the termination) to keV (at the base). The gas emissivity required to produce the X-ray luminosity at level implies hydrogen number densities cm-3 in these regions (assuming a filling factor of unity). Clearly, if the thermal nature of this emission were true, it would have a drastically higher pressure compared to the ambient soft X-ray emitting gas and would expand sideways supersonically, driving a strong shock wave. Alternatively, the emission might result from the population of relativistic particles accelerated within EXJs. The detection of polarization in X-rays (Kaaret et al., 2024) and TeV emission from regions cospatial with the EXJs (Abeysekara et al., 2018; Alfaro et al., 2024; H. E. S. S. Collaboration et al., 2024) strongly supports the non-thermal scenario.
The absence of any noticeable radio, infrared, or optical emission from these structures would then require substantial self-absorption of radiation at these wavelengths (which is unlikely given that the gas number density is likely very low inside them) or a rather peculiar energy distribution of the relativistic particles with a large fraction of them having energies in excess of tens of TeV. Another possibility could be scattering of very bright and beamed emission of the central source (e.g., Panferov & Kabanov, 2013; Khabibullin & Sazonov, 2016, 2019), but the small optical depth with respect to the Thomson scattering, , and the observed spectral gradient along these structures disfavor this possibility.
6 Discussion
The Eastern EXJ was well mapped in X-rays with XMM-Newton, Chandra, and NuSTAR (e.g., Brinkmann et al., 2007; Safi-Harb et al., 2022), revealing its non-thermal nature. Furthermore, recent IXPE observation (Kaaret et al., 2024) revealed polarized X-ray emission from a patch adjacent to the base of the Eastern EXJ. These data suggest that X-ray photons are produced by a synchrotron mechanism with the magnetic field predominantly aligned with the EXJ axis. The eROSITA data show that the bases of both EXJs are located at essentially the same distances from SS 433, have similarly sharp inner boundaries, and non-thermal (featureless) spectra across their entire extents.
These structures have often been considered as a direct consequence of the relativistic jets’ propagation inside the nebula before they deposit the bulk of their kinetic energy in the lobes, energizing them and the whole extent of the nebula (e.g., Panferov, 2014). However, the direction of the SS 433 jets is known to precess about the mean axis with the amplitude degrees, which is a factor of 2 larger than the observed opening angle of EXJ. Since this extended X-ray emission comes from distances of tens of pc, i.e. orders of magnitude further away from the central source than the compact radio and X-ray jets, which are traced up to a few cm, the actual connection between them is not obvious (e.g., Panferov, 2017). It might be governed by recollimation of the jets’ flows at some intermediate distance from the source (Eichler, 1983; Peter & Eichler, 1993), a historical change in the jets’ direction (Goodall et al., 2011a; Monceau-Baroux et al., 2015), or a different regime of baryonic jets propagation through the magnetized ISM after their gas “freezes out” after recombining due to rapid adiabatic expansion, cooling losses and lack of ambient interactions (Churazov et al., 2020).
If we assume that there were no recent drastic changes in the properties of W50/SS 433 (e.g., temporary quenching of the central source), the overall picture of the “large-scale energy flow” in the W50/SS 433, supported by X-ray data, can be schematically decomposed into three major components (illustrated in Fig. 8):
“Dark flow”: from a compact source up to the base of EXJs covering the range of distances from to from SS 433. Over this distance range, there are no visible signatures in the X-ray (or any other) band.
“Non-thermal flow”: from the base to the end of EXJs that have featureless non-thermal spectra (presumably, synchrotron emission).
“Thermal flow”: extended diffuse regions between the W50 radio boundaries and EXJs. Its X-ray emission is typical for ISM shock-heated to sub-keV temperatures.
Analytical and numerical models primarily aimed at reproducing W50’s radio morphology often assume that a quasi-spherical part of the nebula is a remnant of the supernova explosion, while the elongated structures are directly associated with the energy release by the binary system (e.g., Begelman et al., 1980; Zealey et al., 1980; Eichler, 1983; Peter & Eichler, 1993; Velázquez & Raga, 2000; Zavala et al., 2008; Goodall et al., 2011a; Asahina et al., 2014; Monceau-Baroux et al., 2015; Panferov, 2017; Ohmura et al., 2021). In the model considered by Churazov et al. (2024), the entire nebula is powered by the central source. There, an anisotropic outflow/wind from the Hyper-Eddington accretor is assumed - a combination of a more powerful collimated outflow along the orbital axis and a quasi-isotropic wind in all other directions. In this model, the “Dark flow” phase corresponds to the free expansion of the wind, and the onset of the EXJs is associated with the termination shock of the isotropic wind, which initiates shocks in the collimated flow where the pressure rises suddenly. Finally, the thermal part is the shock-heated ISM. In this model, the characteristic shape of the W50 nebula (a quasi-spherical part and two extensions) reflects the anisotropy of the velocity and kinetic power of the wind, with no need for the impact of the narrow baryonic jets at all.
In all these models, the outer shock, which delineates the radio boundary of W50, is not dissimilar to the shocks seen around middle-aged supernova remnants or stellar wind-blown bubbles (e.g., Konigl, 1983; Dubner et al., 1998). What makes W50 “special” is the presence of EXJs, which are radio-faint but bright in X-rays and VHE gamma-rays. These structures are also plausibly related to efficient particle acceleration at strong shocks, but with a shock configuration that is rather different compared to typical SNR shocks (e.g. Bykov et al., 2025). Remarkable symmetry of the EXJs, very clearly demonstrated by the eROSITA image, points towards irrelevance of the ambient ISM properties to the energy and momentum propagation, as well as the particle acceleration regime, meaning that a cocoon or a wind-blown cavity likely encompasses these structures in the current epoch. In this regard, the presence of H emitting filaments in several places across the nebula might help probe the energetic content of the nebula (e.g., Boumis et al., 2007; Abolmasov et al., 2010; Farnes et al., 2017). Interestingly, the location of the brightest H filaments does not correspond to any prominent features in thermal X-ray emission. This might be the case when the H emission comes from the outer boundary of the nebula, where the shock wave is already in the radiative mode. Significantly higher gas temperature obtained for the emission from the Western lobe might also be indicative of this process, namely that the radiative shock wave runs against the Galactic density gradient and stalls earlier in this direction, evacuating a smaller volume to be filled by the hot X-ray gas. Given the presence of bright silicon lines in the spectra extracted from these regions, an observation of them with the XRISM observatory (Tashiro et al., 2025) might provide estimates of the directed and turbulent velocities. A large grasp soft X-ray microcalorimeter mission, like LEM (Kraft et al., 2022) or HUBS (Bregman et al., 2023) would also be perfectly suited for in-depth spectral exploration of this complex system with multiple interacting components.
7 Conclusions
SRG/eROSITA observations of W50/SS 433 provide, for the first time, the entire X-ray map of the nebula with high spatial and spectral resolution. The data support a physical picture in which the anisotropic energy flow from SS 433 evolves passing through three distinct stages:
an invisible “dark” flow between 0.1 and 25 pc (=anisotropic wind from the binary system)
a “non-thermal” flow over another (=EXJs)
a thermal flow (= shock-heated ISM) that envelopes EXJs.
The appearance of the nebula is further affected by the large ambient density gradients and a heavy foreground photoelectric absorption that projects almost exactly on the central quasi-spherical part of the W50 nebula.
The thermal part of the W50 X-ray emission can be reasonably well described by the shock-heated plasma that has not yet reached temperature and ionization equilibrium. Such emission is typical for middle-age or old SNRs. The outer radio boundary of the nebula is also reminiscent of the SNR shocks.
On the contrary, the “Extended X-ray Jets” are the most remarkable features of this system on scales of tens of pc. Their sharp inner edges plausibly correspond to extreme shocks that accelerate particles and power the X-ray (synchrotron) emission, and, at 10 orders of magnitude higher energies, the TeV emission. W50/SS 433 system illustrates clearly the important role the Hyper-Eddington accretors might play for the energetics of the ISM in galaxies at different redshifts and the production of ultra-high-energy particles.
Acknowledgements.
This work is partly based on observations with the eROSITA telescope onboard SRG space observatory. The SRG observatory was built by Roskosmos in the interests of the Russian Academy of Sciences, represented by its Space Research Institute (IKI) in the framework of the Russian Federal Space Program, with the participation of the Deutsches Zentrum für Luft- und Raumfahrt (DLR). The eROSITA X-ray telescope was built by a consortium of German Institutes led by MPE, and supported by DLR. The SRG spacecraft was designed, built, launched, and operated by the Lavochkin Association and its subcontractors. The science data are downlinked via the Deep Space Network Antennae in Bear Lakes, Ussurijsk, and Baikonur, funded by Roskosmos. The development and construction of the eROSITA X-ray instrument was led by MPE, with contributions from the Dr. Karl Remeis Observatory Bamberg ECAP (FAU Erlangen-Nuernberg), the University of Hamburg Observatory, the Leibniz Institute for Astrophysics Potsdam (AIP), and the Institute for Astronomy and Astrophysics of the University of Tübingen, with the support of DLR and the Max Planck Society. The Argelander Institute for Astronomy of the University of Bonn and the Ludwig Maximilians Universität München also participated in the science preparation for eROSITA. The eROSITA data were processed using the eSASS/NRTA software system developed by the German eROSITA consortium and analyzed using proprietary data reduction software developed by the Russian eROSITA Consortium. IK acknowledges support by the COMPLEX project from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program grant agreement ERC-2019-AdG 882679.References
- Abell & Margon (1979) Abell, G. O. & Margon, B. 1979, Nature, 279, 701
- Abeysekara et al. (2018) Abeysekara, A. U., Albert, A., Alfaro, R., et al. 2018, Nature, 562, 82
- Abolmasov et al. (2010) Abolmasov, P., Maryeva, O., & Burenkov, A. N. 2010, AN, 331, 412
- Alfaro et al. (2024) Alfaro, R., Alvarez, C., Arteaga-Velázquez, J. C., et al. 2024, ApJ, 976, 30
- Arnason et al. (2021) Arnason, R. M., Papei, H., Barmby, P., Bahramian, A., & Gorski, M. D. 2021, MNRAS, 502, 5455
- Arnaud (1996) Arnaud, K. A. 1996, in Astronomical Society of the Pacific Conference Series, Vol. 101, Astronomical Data Analysis Software and Systems V, ed. G. H. Jacoby & J. Barnes, 17
- Asahina et al. (2014) Asahina, Y., Ogawa, T., Kawashima, T., et al. 2014, ApJ, 789, 79
- Atapin & Fabrika (2016) Atapin, K. E. & Fabrika, S. N. 2016, AstL, 42, 517
- Begelman et al. (2006) Begelman, M. C., King, A. R., & Pringle, J. E. 2006, MNRAS, 370, 399
- Begelman et al. (1980) Begelman, M. C., Sarazin, C. L., Hatchett, S. P., McKee, C. F., & Arons, J. 1980, ApJ, 238, 722
- Blundell & Bowler (2004) Blundell, K. M. & Bowler, M. G. 2004, ApJL, 616, L159
- Blundell et al. (2011) Blundell, K. M., Schmidtobreick, L., & Trushkin, S. 2011, MNRAS, 417, 2401
- Borkowski et al. (2001) Borkowski, K. J., Lyerly, W. J., & Reynolds, S. P. 2001, ApJ, 548, 820
- Boumis et al. (2007) Boumis, P., Meaburn, J., Alikakos, J., et al. 2007, MNRAS, 381, 308
- Bregman et al. (2023) Bregman, J., Cen, R., Chen, Y., et al. 2023, SCPMA, 66, 299513
- Brinkmann et al. (1996) Brinkmann, W., Aschenbach, B., & Kawai, N. 1996, A&A, 312, 306
- Brinkmann et al. (1988) Brinkmann, W., Fink, H. H., Massaglia, S., Bodo, G., & Ferrari, A. 1988, A&A, 196, 313
- Brinkmann & Kawai (2000) Brinkmann, W. & Kawai, N. 2000, A&A, 363, 640
- Brinkmann et al. (1991) Brinkmann, W., Kawai, N., Matsuoka, M., & Fink, H. H. 1991, A&A, 241, 112
- Brinkmann et al. (2005) Brinkmann, W., Kotani, T., & Kawai, N. 2005, A&A, 431, 575
- Brinkmann et al. (2007) Brinkmann, W., Pratt, G. W., Rohr, S., Kawai, N., & Burwitz, V. 2007, A&A, 463, 611
- Broderick et al. (2018) Broderick, J. W., Fender, R. P., Miller-Jones, J. C. A., et al. 2018, MNRAS, 475, 5360
- Bykov et al. (2025) Bykov, A. M., Osipov, S. M., Romansky, V. I., et al. 2025, PhRvD, 112, 063017
- Cherepashchuk et al. (2023) Cherepashchuk, A., Belinski, A., Dodin, A., & Postnov, K. 2023, NewA, 103, 102060
- Cherepashchuk (1981) Cherepashchuk, A. M. 1981, MNRAS, 194, 761
- Cherepashchuk et al. (2025) Cherepashchuk, A. M., Dodin, A. V., & Postnov, K. A. 2025, arXiv, arXiv:2506.01106
- Cherepashchuk et al. (2018) Cherepashchuk, A. M., Esipov, V. F., Dodin, A. V., Davydov, V. V., & Belinskii, A. A. 2018, ARep, 62, 747
- Chi et al. (2024) Chi, Y.-H., Huang, J., Zhou, P., et al. 2024, ApJL, 975, L28
- Churazov et al. (2023) Churazov, E., Khabibullin, I., Bykov, A. M., Lyskova, N., & Sunyaev, R. 2023, A&A, 670, A156
- Churazov et al. (2021a) Churazov, E., Khabibullin, I., Lyskova, N., Sunyaev, R., & Bykov, A. M. 2021a, A&A, 651, A41
- Churazov et al. (2025) Churazov, E., Khabibullin, I., Lyskova, N., Sunyaev, R., & Dolag, K. 2025, arXiv, arXiv:2507.19987
- Churazov et al. (2020) Churazov, E., Khabibullin, I., & Sunyaev, R. 2020, MNRAS, 495, L51
- Churazov et al. (2024) Churazov, E. M., Khabibullin, I. I., & Bykov, A. M. 2024, A&A, 688, A4
- Churazov et al. (2021b) Churazov, E. M., Khabibullin, I. I., Bykov, A. M., et al. 2021b, MNRAS, 507, 971
- Dubner et al. (1998) Dubner, G. M., Holdaway, M., Goss, W. M., & Mirabel, I. F. 1998, AJ, 116, 1842
- Eichler (1983) Eichler, D. 1983, ApJ, 272, 48
- Eikenberry et al. (2001) Eikenberry, S. S., Cameron, P. B., Fierce, B. W., et al. 2001, ApJ, 561, 1027
- Elston & Baum (1987) Elston, R. & Baum, S. 1987, AJ, 94, 1633
- Fabian & Rees (1979) Fabian, A. C. & Rees, M. J. 1979, MNRAS, 187, 13P
- Fabrika (2004) Fabrika, S. 2004, ASPRv, 12, 1
- Fabrika & Mescheryakov (2001) Fabrika, S. & Mescheryakov, A. 2001, in IAU Symposium, Vol. 205, Galaxies and their Constituents at the Highest Angular Resolutions, ed. R. T. Schilizzi, 268
- Farnes et al. (2017) Farnes, J. S., Gaensler, B. M., Purcell, C., et al. 2017, MNRAS, 467, 4777
- Filippova et al. (2006) Filippova, E., Revnivtsev, M., Fabrika, S., Postnov, K., & Seifina, E. 2006, A&A, 460, 125
- Fogantini et al. (2023) Fogantini, F. A., García, F., Combi, J. A., et al. 2023, A&A, 669, A149
- Forman et al. (1978) Forman, W., Jones, C., Cominsky, L., et al. 1978, ApJS, 38, 357
- Fragile et al. (2025) Fragile, P. C., Middleton, M. J., Bollimpalli, D. A., & Smith, Z. 2025, MNRAS, 540, 2820
- Fuchs et al. (2006) Fuchs, Y., Koch Miramond, L., & Ábrahám, P. 2006, A&A, 445, 1041
- Gaia Collaboration et al. (2023) Gaia Collaboration, Vallenari, A., Brown, A. G. A., et al. 2023, A&A, 674, A1
- Geldzahler et al. (1980) Geldzahler, B. J., Pauls, T., & Salter, C. J. 1980, A&A, 84, 237
- Goodall et al. (2011a) Goodall, P. T., Alouani-Bibi, F., & Blundell, K. M. 2011a, MNRAS, 414, 2838
- Goodall et al. (2011b) Goodall, P. T., Blundell, K. M., & Bell Burnell, S. J. 2011b, MNRAS, 414, 2828
- Goranskij (2011) Goranskij, V. 2011, PZ, 31, 5
- Green (2025) Green, D. A. 2025, JApA, 46, 14
- Green et al. (2019) Green, G. M., Schlafly, E., Zucker, C., Speagle, J. S., & Finkbeiner, D. 2019, ApJ, 887, 93
- Gregory et al. (1996) Gregory, P. C., Scott, W. K., Douglas, K., & Condon, J. J. 1996, ApJS, 103, 427
- Grindlay et al. (1984) Grindlay, J. E., Band, D., Seward, F., et al. 1984, ApJ, 277, 286
- H. E. S. S. Collaboration et al. (2024) H. E. S. S. Collaboration, Aharonian, F., Ait Benkhali, F., et al. 2024, Sci, 383, 402
- Heinz & Sunyaev (2002) Heinz, S. & Sunyaev, R. 2002, A&A, 390, 751
- Hjellming & Johnston (1981a) Hjellming, R. M. & Johnston, K. J. 1981a, ApJL, 246, L141
- Hjellming & Johnston (1981b) Hjellming, R. M. & Johnston, K. J. 1981b, Nature, 290, 100
- Holden & Caswell (1969) Holden, D. J. & Caswell, J. L. 1969, MNRAS, 143, 407
- Jiang et al. (2019) Jiang, Y.-F., Stone, J. M., & Davis, S. W. 2019, ApJ, 880, 67
- Kaaret et al. (2017) Kaaret, P., Feng, H., & Roberts, T. P. 2017, ARA&A, 55, 303
- Kaaret et al. (2024) Kaaret, P., Ferrazzoli, R., Silvestri, S., et al. 2024, ApJL, 961, L12
- Katz (1980) Katz, J. I. 1980, ApJL, 236, L127
- Katz et al. (1982) Katz, J. I., Anderson, S. F., Margon, B., & Grandi, S. A. 1982, ApJ, 260, 780
- Kawai et al. (1989) Kawai, N., Matsuoka, M., Pan, H.-C., & Stewart, G. C. 1989, PASJ, 41, 491
- Khabibullin et al. (2022) Khabibullin, I., Churazov, E., & Sunyaev, R. 2022, MNRAS, 509, 6068
- Khabibullin et al. (2016) Khabibullin, I., Medvedev, P., & Sazonov, S. 2016, MNRAS, 455, 1414
- Khabibullin & Sazonov (2016) Khabibullin, I. & Sazonov, S. 2016, MNRAS, 457, 3963
- Khabibullin et al. (2023) Khabibullin, I. I., Churazov, E. M., Bykov, A. M., Chugai, N. N., & Sunyaev, R. A. 2023, MNRAS, 521, 5536
- Khabibullin et al. (2024) Khabibullin, I. I., Churazov, E. M., Chugai, N. N., et al. 2024, A&A, 689, A278
- Khabibullin & Sazonov (2012) Khabibullin, I. I. & Sazonov, S. Y. 2012, AstL, 38, 443
- Khabibullin & Sazonov (2017) Khabibullin, I. I. & Sazonov, S. Y. 2017, AstL, 43, 388
- Khabibullin & Sazonov (2019) Khabibullin, I. I. & Sazonov, S. Y. 2019, AstL, 45, 282
- Kimura et al. (2020) Kimura, S. S., Murase, K., & Mészáros, P. 2020, ApJ, 904, 188
- Kirshner & Chevalier (1980) Kirshner, R. P. & Chevalier, R. A. 1980, ApJL, 242, L77
- Konigl (1983) Konigl, A. 1983, MNRAS, 205, 471
- Kotani et al. (1996) Kotani, T., Kawai, N., Matsuoka, M., & Brinkmann, W. 1996, PASJ, 48, 619
- Kraft et al. (2022) Kraft, R., Markevitch, M., Kilbourne, C., et al. 2022, arXiv, arXiv:2211.09827
- Kubota et al. (2010) Kubota, K., Ueda, Y., Kawai, N., et al. 2010, PASJ, 62, 323
- Leibowitz (1984) Leibowitz, E. M. 1984, MNRAS, 210, 279
- LHAASO Collaboration (2024) LHAASO Collaboration. 2024, arXiv, arXiv:2410.08988
- Lopez et al. (2006) Lopez, L. A., Marshall, H. L., Canizares, C. R., Schulz, N. S., & Kane, J. F. 2006, ApJ, 650, 338
- Margon (1984) Margon, B. 1984, ARA&A, 22, 507
- Margon et al. (1979) Margon, B., Ford, H. C., Katz, J. I., et al. 1979, ApJL, 230, L41
- Marshall et al. (2013) Marshall, H. L., Canizares, C. R., Hillwig, T., et al. 2013, ApJ, 775, 75
- Marshall et al. (2002) Marshall, H. L., Canizares, C. R., & Schulz, N. S. 2002, ApJ, 564, 941
- Medvedev & Fabrika (2010) Medvedev, A. & Fabrika, S. 2010, MNRAS, 402, 479
- Medvedev et al. (2019) Medvedev, P. S., Khabibullin, I. I., & Sazonov, S. Y. 2019, AstL, 45, 299
- Medvedev et al. (2018) Medvedev, P. S., Khabibullin, I. I., Sazonov, S. Y., Churazov, E. M., & Tsygankov, S. S. 2018, AstL, 44, 390
- Medvedev et al. (2022) Medvedev, P. S., Khabibullin, I. I., Semena, A. N., et al. 2022, AstL, 48, 389
- Middleton et al. (2021) Middleton, M. J., Walton, D. J., Alston, W., et al. 2021, MNRAS, 506, 1045
- Migliari et al. (2005) Migliari, S., Fender, R. P., Blundell, K. M., Méndez, M., & van der Klis, M. 2005, MNRAS, 358, 860
- Miller-Jones et al. (2008) Miller-Jones, J. C. A., Migliari, S., Fender, R. P., et al. 2008, ApJ, 682, 1141
- Monceau-Baroux et al. (2015) Monceau-Baroux, R., Porth, O., Meliani, Z., & Keppens, R. 2015, A&A, 574, A143
- Namiki et al. (2000) Namiki, M., Kawai, N., Kotani, T., Mamauchi, S., & Brinkmann, W. 2000, AdSpR, 25, 709
- Ohmura et al. (2021) Ohmura, T., Ono, K., Sakemi, H., et al. 2021, ApJ, 910, 149
- Ohsuga & Mineshige (2011) Ohsuga, K. & Mineshige, S. 2011, ApJ, 736, 2
- Panferov (2014) Panferov, A. 2014, A&A, 562, A130
- Panferov (2017) Panferov, A. A. 2017, A&A, 599, A77
- Panferov & Kabanov (2013) Panferov, A. A. & Kabanov, A. A. 2013, arXiv, arXiv:1306.4486
- Peretti et al. (2025) Peretti, E., Petropoulou, M., Vasilopoulos, G., & Gabici, S. 2025, A&A, 698, A188
- Peter & Eichler (1993) Peter, W. & Eichler, D. 1993, ApJ, 417, 170
- Poutanen et al. (2007) Poutanen, J., Lipunova, G., Fabrika, S., Butkevich, A. G., & Abolmasov, P. 2007, MNRAS, 377, 1187
- Predehl et al. (2021) Predehl, P., Andritschke, R., Arefiev, V., et al. 2021, A&A, 647, A1
- Predehl et al. (2020) Predehl, P., Sunyaev, R. A., Becker, W., et al. 2020, Nature, 588, 227
- Rosado et al. (2021) Rosado, M., Sánchez-Cruces, M., Ambrocio-Cruz, P., & Trejo, D. 2021, MNRAS, 506, 4263
- Safi-Harb et al. (2022) Safi-Harb, S., Mac Intyre, B., Zhang, S., et al. 2022, ApJ, 935, 163
- Safi-Harb & Ögelman (1997) Safi-Harb, S. & Ögelman, H. 1997, ApJ, 483, 868
- Safi-Harb & Petre (1999) Safi-Harb, S. & Petre, R. 1999, ApJ, 512, 784
- Sakai et al. (2025) Sakai, Y., Yamada, S., Sakemi, H., et al. 2025, arXiv, arXiv:2507.19042
- Sakemi et al. (2021) Sakemi, H., Omae, R., Ohmura, T., & Machida, M. 2021, PASJ, 73, 530
- Sarazin et al. (1980) Sarazin, C. L., Begelman, M. C., & Hatchett, S. P. 1980, ApJL, 238, L129
- Seward et al. (1980) Seward, F., Grindlay, J., Seaquist, E., & Gilmore, W. 1980, Nature, 287, 806
- Seward et al. (1976) Seward, F. D., Page, C. G., Turner, M. J. L., & Pounds, K. A. 1976, MNRAS, 175, 39P
- Shakura & Sunyaev (1973) Shakura, N. I. & Sunyaev, R. A. 1973, A&A, 24, 337
- Sharpless (1959) Sharpless, S. 1959, ApJS, 4, 257
- Shklovskii (1981) Shklovskii, I. S. 1981, SvA, 25, 315
- Shuder et al. (1980) Shuder, J. M., Hatfield, B. F., & Cohen, R. D. 1980, PASP, 92, 259
- Spencer (1979) Spencer, R. E. 1979, Nature, 282, 483
- Stephenson & Sanduleak (1977) Stephenson, C. B. & Sanduleak, N. 1977, ApJS, 33, 459
- Stewart et al. (1987) Stewart, G. C., Watson, M. G., Matsuoka, M., et al. 1987, MNRAS, 228, 293
- Sunyaev et al. (2021) Sunyaev, R., Arefiev, V., Babyshkin, V., et al. 2021, A&A, 656, A132
- Tashiro et al. (2025) Tashiro, M., Kelley, R., Watanabe, S., et al. 2025, PASJ
- Tsuji et al. (2025) Tsuji, N., Inoue, Y., Khangulyan, D., et al. 2025, arXiv, arXiv:2510.06431
- van den Heuvel (1981) van den Heuvel, E. P. J. 1981, VA, 25, 95
- van den Heuvel et al. (1980) van den Heuvel, E. P. J., Ostriker, J. P., & Petterson, J. A. 1980, A&A, 81, L7
- van den Heuvel et al. (2017) van den Heuvel, E. P. J., Portegies Zwart, S. F., & de Mink, S. E. 2017, MNRAS, 471, 4256
- Velázquez & Raga (2000) Velázquez, P. F. & Raga, A. C. 2000, A&A, 362, 780
- Wang et al. (2025) Wang, J., Reville, B., & Aharonian, F. A. 2025, ApJL, 989, L25
- Watson et al. (1986) Watson, M. G., Stewart, G. C., Brinkmann, W., & King, A. R. 1986, MNRAS, 222, 261
- Watson et al. (1983) Watson, M. G., Willingale, R., Grindlay, J. E., & Seward, F. D. 1983, ApJ, 273, 688
- Westerhout (1958) Westerhout, G. 1958, BAN, 14, 215
- Whitmire & Matese (1980) Whitmire, D. P. & Matese, J. J. 1980, MNRAS, 193, 707
- Wilms et al. (2000) Wilms, J., Allen, A., & McCray, R. 2000, ApJ, 542, 914
- Yamauchi et al. (1994) Yamauchi, S., Kawai, N., & Aoki, T. 1994, PASJ, 46, L109
- Yoshioka et al. (2022) Yoshioka, S., Mineshige, S., Ohsuga, K., Kawashima, T., & Kitaki, T. 2022, PASJ, 74, 1378
- Zavala et al. (2008) Zavala, J., Velázquez, P. F., Cerqueira, A. H., & Dubner, G. M. 2008, MNRAS, 387, 839
- Zealey et al. (1980) Zealey, W. J., Dopita, M. A., & Malin, D. F. 1980, MNRAS, 192, 731
Appendix A Spectrum of the central source
As one can see in Fig. 6, both blue-shifted and red-shifted lines of highly ionized (H-like and He-like) heavy elements (in particular, silicon, sulfur, iron, and nickel) moving at 0.26 speed of light can be easily discriminated and identified. They arise from the approaching (with apparent redshift -0.085) and the receding jet (with apparent redshift 0.167), respectively. In agreement with the findings of numerous previous observations (see Medvedev et al. 2019, for a recent summary), a factor of 10 overabundance of nickel (with respect to iron) in the gas of the jets is required to describe the spectrum well (most clearly indicated by the ratio of the Ni XXVII and Fe XXV lines). Spectral resolution of eROSITA is not sufficient to resolve the width of the lines, which is known to be dominated by the ballistic motion of gas within a cone of with opening angle , as measured by high resolution spectroscopy with Chandra gratings (e.g. Marshall et al. 2002). Some residual emission can also be observed in the iron K complex region between 6 and 7 keV, which could come from the reflection of central source emission on the walls of the optically thick accretion disk funnel (Medvedev & Fabrika 2010) or in the optically thin accretion disk wind (Medvedev et al. 2018, 2019). Comparison of the spectra taken in the two epochs of observations separated by one orbital period shows very good consistency between them333Since one orbital period separation is small compared to the precession period and the observations were close to an extreme point of the precession curve, the change in the Doppler shifts of the jets is small., with the biggest difference observed in the 6-7 keV band. This might be related to the change in the visible geometry of this region or to the physical conditions within it (see a discussion in Medvedev et al. 2019). High-resolution spectroscopy of this emission with the XRISM observatory (Tashiro et al. 2025) will open new possibilities for characterizing this emission and potentially give access to more intricate diagnostics based on velocity structure and resonant scattering in the jets (Khabibullin & Sazonov 2012), as well as possible signatures of multiphase cooling and ionization state dynamics in supersonically expanding jets prone to thermal instability and fragmentation (Brinkmann et al. 1988; Brinkmann & Kawai 2000; Khabibullin et al. 2016). The SS433’s spectrum presented here illustrates the spectroscopic capabilities of eROSITA and shows that, for sufficiently bright sources, reliable continuum and line characterization can be achieved up to 10 keV.
Appendix B Parameters of the spectral models
Here we list parameters of the best-fit models used to describe spectra shown in Fig. 7 and discussed in Sec. 5. The spectral analysis was performed using standard routines of the XSPEC package (Arnaud 1996). Namely, the absorbed (using tbabs model by Wilms et al. 2000) non-equilibrium plasma emission model (nei model by Borkowski et al. 2001) was used to describe thermal emission, while similarly absorbed thermal bremsstrahlung model (tbabs*brems in XSPEC notation) was used for a phenomenological description of the non-thermal emission from the EXJs. Due to the complexity of the system, the X-ray emission from all these regions is likely a mixture of several components; hence, one needs to take these derived parameters with caution and probably as approximate guidelines for the physical conditions within them. For the background estimation, adjacent regions of significant size were used, and the signal inside them was subtracted from the corresponding source region. For all regions except for the relatively faint Regions 4 and 8, the relative contribution of the background emission in the source regions was rather small.
| # | Norm | |||
| keV | ||||
| Thermal regions (tbabs*vnei) | ||||
| 1 | 0.670.02 | 0.290.02 | 0.260.05 | 12.333.63 |
| 3 | 0.560.02 | 0.560.07 | 0.110.01 | 0.800.19 |
| 4 | 0.790.05 | 0.950.25 | 0.090.01 | 0.210.07 |
| 6 | 0.850.03 | 0.760.06 | 0.640.12 | 2.010.28 |
| 7 | 0.720.02 | 1.300.18 | 0.080.01 | 0.880.13 |
| 8 | 0.750.05 | 0.780.22 | 0.130.05 | 0.420.18 |
| Non-thermal regions (tbabs*brems) | ||||
| 2 | 0.510.01 | 3.050.15 | - | 1.620.04 |
| 5 | 0.840.03 | 14.53.9 | - | 0.640.01 |
Appendix C Impact of foreground absorption
In this section, we demonstrate that some of the large-scale variations in the diffuse X-ray emission from W50 is plausibly caused by foreground absorption. To do so, we used the 3D dust maps from Bayestar19 (see Green et al. 2019, for detailed description). We are looking for the intervening absorption between W50 and us, i.e., physically unrelated layers of the absorbing material. Given the adopted distance to SS 433 of 5 kpc, we integrated extinction () only up to a distance of 3 kpc from the Sun. The resulting extinction map is superposed on the X-ray image in Fig. 9. A prominent wide dust lane is seen in this image that crosses the central part of the W50 nebula diagonally from NE to SW. Typical values of the extinction in this region are , which translate into equivalent hydrogen column densities . Therefore, soft X-ray emission from the central part of W50 can be heavily absorbed (several orders of magnitude at 1 keV). While some geometric correspondences between the X-ray and extinction maps look tantalizing, we conclude that most probably this is a by-chance projection. Therefore, the lack of observed soft X-ray emission from the quasi-circular part of the nebula does not place significant constraints on the properties of W50. The soft X-ray emission might be there, but we can not see it.